# FactSelfCheck: Fact-Level Black-Box Hallucination Detection for LLMs

### v2

@@ Abstract
Large Language Models (LLMs) frequently generate hallucinated content, posing significant challenges for applications where factuality is crucial. While existing hallucination detection methods typically operate at the sentence level or passage level, we propose FactSelfCheck, a novel black-box sampling-based method that enables fine-grained fact-level detection. Our approach represents text as knowledge graphs consisting of facts in the form of triples. Through analyzing factual consistency across multiple LLM responses, we compute fine-grained hallucination scores without requiring external resources or training data. Our evaluation demonstrates that FactSelfCheck performs competitively with leading sentence-level sampling-based methods while providing more detailed insights. Most notably, our fact-level approach significantly improves hallucination correction, achieving a $35.5\%$ increase in factual content compared to the baseline, while sentence-level SelfCheckGPT yields only a $10.6\%$ improvement. The granular nature of our detection enables more precise identification and correction of hallucinated content. Additionally, we contribute a new dataset for evaluating sampling-based methods – FavaMultiSamples.
@@ 1 Introduction
Large Language Models (LLMs) have gained significant attention from academia and industry recently. However, a major limitation of LLMs is their tendency to generate hallucinated information Farquhar et al. (2024); Huang et al. (2025), posing significant challenges for applications where factual correctness is crucial, such as healthcare Sallam (2023). Although numerous methods have been proposed to reduce hallucinations Zhang et al. (2023), it is not possible to eliminate them, and LLMs will constantly hallucinate Lee (2023); Xu et al. (2024). Therefore, there remains a critical need for reliable hallucination detection in LLM responses. Effective detection enables system interventions by either preventing the transmission of hallucinated content to users or facilitating its correction Zhang et al. (2023).
Previous approaches to hallucination detection have primarily focused on classifying hallucinations at either the passage or sentence level Huang et al. (2025). While valuable, these approaches are limited in their granularity, as they do not provide detailed information about specific hallucinated facts. To address this limitation, we propose a novel method for hallucination detection that operates at the fact level, offering finer-grained analysis. In our approach, we define a fact as a triple consisting of a head, relation, and tail – a standard representation in knowledge graphs (e.g., (Robert Smithmember ofThe Cure)) Hamilton et al. (2017). Our method provides more precise and actionable information by computing hallucination scores for individual facts than traditional passage-level or sentence-level classification approaches. Moreover, our method represents facts in a structured way through knowledge graphs (KGs), enabling more straightforward interpretation, processing, and verification of factual content Pan et al. (2024).
@@
Our granular approach is motivated by two key observations. First, a single sentence can contain multiple facts, with the number of facts varying significantly across sentences, contexts, and domains. This variability makes it challenging to identify hallucinated aspects of generated output precisely when using sentence-level detection. Second, false information can be dispersed throughout a text, as a single fact may appear across multiple sentences. That can mislead the sentence-level detection, as the factuality of a sentence is dependent on the previous sentences Zhang et al. (2024a). These challenges become particularly acute when analyzing long texts generated by LLMs. Fine-grained fact-level detection provides a more precise understanding of text factuality than a sentence analysis. It enables better assessment of content reliability and, as we show later, more effective factuality correction.
@@
We propose FactSelfCheck, a black-box method for fact-level hallucination detection, meaning it does not require access to the model’s internal parameters. This design choice makes our approach universally applicable across any LLM, including closed models, like GPT OpenAI et al. (2024). Following a sampling-based detection paradigm, introduced by Manakul et al., 2023, our method utilizes multiple response generations and analyzes the factual consistency of extracted facts across these samples. This paradigm is based on the phenomenon that factual information remains largely consistent across different generations, while hallucinated content tends to vary or contradict itself between samples Manakul et al. (2023); Wang et al. (2023). This way, we can effectively identify hallucinated facts without relying on external resources (zero-resource) or access to the model’s internal parameters (black-box). Moreover, our method is non-parametric, as it does not require any training, making it easy to implement in any domain without the need for training data. The FactSelfCheck pipeline consists of three main steps: knowledge graph extraction, which extracts sets of facts from the initial response and samples; fact-level hallucination scoring; and calculating sentence-level scores by aggregating fact-level scores.
@@
We evaluated our method using the WikiBio GPT-3 Hallucination Dataset Manakul et al. (2023) and the FavaMultiSamples dataset, a new dataset for evaluating sampling-based methods. We performed sentence-level evaluation by aggregating fact-level scores. Our approach achieves performance comparable to leading sampling-based methods while providing more detailed information about hallucinations. Additionally, we demonstrate that our fact-level approach significantly improves hallucination correction. Compared to a baseline, providing incorrect facts to the correction method leads to a $35.5\%$ increase in factual content, while passing incorrect sentences leads to an $10.6\%$ increase.
@@
    The novel black-box sampling-based method for fact-level hallucination detection – FactSelfCheck. It enables fine-grained hallucination detection in LLM responses without requiring training data or external resources, as it is both non-parametric and zero-resource. We propose two effective approaches for measuring factual consistency across multiple samples: FactSelfCheck-KG using knowledge graph comparisons and FactSelfCheck-Text using direct text comparison.
@@ 2 Related work
Xu et al., 2024 have proven that hallucinations are inevitable in LLMs. As LLMs are powerful tools, many recent studies have been conducted regarding hallucination mitigation and detection Zhang et al. (2023); Huang et al. (2025). The detection methods can be divided into two groups: white-box and black-box.
@@
Black-box approaches operate without access to the model’s internal states and aim to detect hallucinations based solely on the text generated by LLMs. Some of these methods use external resources to collect evidence (Min et al., 2023; Chern et al., 2023). Others leverage LLMs to detect hallucinations like CoVe Dhuliawala et al. (2024), which utilizes the chain-of-thought paradigm for detection. Another category is sampling-based methods, such as SelfCheckGPT (Manakul et al., 2023), which evaluate factuality by generating multiple responses (stochastic samples) and assessing consistency between the original response and these samples. The paradigm of utilizing LLM to check it’s own responses was widely studied and adopted in many works (Kadavath et al., 2022; Lin et al., 2024; Ferraz et al., 2024; Zhang et al., 2024b; Miao et al., 2023). Many of these approaches employ a multi-step decomposition strategy to break down the complex task of hallucination detection into more manageable subtasks, a methodology we also adopt in our approach.
The most popular approach is to classify hallucinations at sentence-level or passage-level Huang et al. (2025). Few methods have been specifically designed to detect hallucinations at the fact level (where facts are defined as triples). GraphEval Sansford et al. (2024) generates a KG from LLM output and compares it with the context provided in the LLM input. FactAlign Rashad et al. (2024) builds KGs from input and output, then compares them after performing entity alignment, a technique that pairs the same entity in different KGs. Most similar to our approach is GCA Fang et al. (2024), which constructs KGs from the response and samples and then compares them by aggregating multiple scores. GCA has significant methodological concerns - they tuned several parameters directly on the evaluation set (which is the only set in this dataset), thus we do not compare against it despite using the same dataset. Moreover, our method belongs to a fundamentally different class as it is zero-shot, requiring no parameter tuning.
@@ 3 Method
We propose FactSelfCheck, a black-box sampling-based method for fact-level hallucination detection, as illustrated in Figure 1.
@@ 3.3 Knowledge Graph Extraction
We adopt an approach that decomposes the knowledge graph extraction task into simpler subtasks, similarly to Edge et al., 2024. This process involves three primary steps: extracting entities, identifying relations, and formulating facts, which are implemented as a sequence of LLM prompts.
@@ 3.4.1 FactSelfCheck-KG
where $\mathcal{H}_{\text{fact}}(f)$ is the hallucination score for fact $f$, and $\mathbb{I}\{f\in KG_{s_{n}}\}$ is an indicator function that equals $1$ if fact $f$ appears in $KG_{s_{n}}$ and 00 otherwise. The higher the $\mathcal{H}_{\text{fact}}$ value, the higher the plausibility of hallucination.
@@
$\Psi(\cdot)=\begin{cases}0&\text{if the LLM returns 'yes'}\\ 1&\text{if the LLM returns 'no'}\end{cases}$ (9)
@@ 3.5 Sentence-Level Hallucination Score
While detecting hallucinations at the fact level offers fine-grained insights, there are scenarios where sentence-level detection is necessary. To achieve this, we aggregate fact-level scores to compute sentence-level scores $\mathcal{H}_{\text{sentence}}(u)$.
@@
where $u$ represents a single sentence, and $U$ is the set of sentences of the response $p$. The aggregation function Agg can be based on operations such as $mean$ or $max$. The $mean$ function provides a smoothed factuality score, whereas $max$ is sensitive to the most hallucinated facts or sentences.
@@ 4 Experimental Setup
In this section, we describe the experimental setup, including the used data, implementation details, and aspects we investigated. Additionally, we conducted a fact-level evaluation to provide a more comprehensive analysis. The detailed description of it is provided in Appendix C.
@@ 4.1 Evaluation Data
Since the method is designed for detecting hallucinations in long generated passages, finding appropriate datasets was challenging. We evaluated our method two datasets. The first one is the WikiBio GPT-3 Hallucination Dataset Manakul et al. (2023) 444huggingface.co/datasets/potsawee/wiki_bio_gpt3_hallucination (cc-by-sa-3.0 license) (later referred to as WikiBio). To the best of our knowledge, this was the only dataset specifically designed for evaluating sampling-based hallucination detection methods. However, the limited availability of such datasets could hinder the development and evaluation of these approaches. To address this limitation, we developed FavaMultiSamples, a novel dataset specifically designed for evaluating methods that analyze multiple samples. We built it upon the FAVA dataset Mishra et al. (2024) and the detailed description of creation is provided in Appendix A. The detailed statistics of the datasets are provided in Appendix B.
Both datasets contain only test data, making it methodologically incorrect to tune parameters, including classification thresholds, on these datasets. While the datasets focus on sentence-level evaluation, our approach provides more fine-grained insights through fact-level analysis. To ensure a meaningful comparison with SelfCheckGPT, we evaluated these levels using aggregation approaches (see Section 3.5). For WikiBio, we followed the protocol established by Manakul et al., 2023, merging the labels major-inaccurate and minor-inaccurate into a single hallucination class.
@@ 4.2 Baseline Models
[…]
We include the two best-performing variants of SelfCheckGPT: Prompt and NLI. We implemented the Prompt variant using the same LLM employed in our method, namely Llama-3.1-70B-Instruct, whereas the original method utilized an unspecified release of GPT-3.5-turbo. For the NLI variant, we used the same model as was used in the original paper555huggingface.co/potsawee/deberta-v3-large-mnli license: Apache license 2.0. Finally, the AttentionScore leverages attention maps (Sriramanan et al., 2024), which is, to the best of our knowledge, the only unsupervised internal state-based method. Its unsupervised nature was crucial since the available datasets do not provide training data. We adapted the AttentionScore for sentence-level detection by implementing two variants: (1) Absolute, which analyzes the complete attention map from the input start to the end of the given sentence, and (2) Relative, which analyzes on the attention map between the start and end of the given sentence.
@@ 4.3 Implementation Details
We employed the Llama-3.1-70B-Instruct model Grattafiori et al. (2024) as the LLM in all steps of our method and for baselines requiring access to LLM. We hosted it locally using vLLM Kwon et al. (2023) on a server with 2xNvidia H100 94GB. For hallucination correction, we utilized GPT-4o OpenAI et al. (2024), as motivated in Section 4.5. We set the LLM’s temperature to $0.0$ for all calls, except during hallucination correction (see Section 4.5), where we set it to $0.5$. The used prompts are available in the code repository. We implemented the methods and experiments using LangChain Chase (2022), Hugging Face Transformers Wolf et al. (2020) and Hugging Face Datasets Lhoest et al. (2021). All pipeline steps were defined using DVC Kuprieiev et al. (2025) to facilitate reproducibility.
To determine whether a language model’s response was ’yes’ or ’no’ in Equation 9, we parsed the text into individual words and verified the presence of the words ’yes’ or ’no’. If either word was detected, the response was excluded from the averaging process in Equations 8 and 10. Moreover, our pipeline is vulnerable to not detecting facts in short, uninformative sentences; for these cases, we set the score to $0.5$. The analysis of the number of such sentences is provided in Appendix D.3.
4.4 Evaluation of Sentence-Level Hallucination Detection

### v3

@@ Abstract
Large Language Models (LLMs) frequently generate hallucinated content, posing significant challenges for applications where factuality is crucial. While existing hallucination detection methods typically operate at the sentence level or passage level, we propose FactSelfCheck, a novel zero-resource black-box sampling-based method that enables fine-grained fact-level detection. Our approach represents long-form text as interpretable knowledge graphs consisting of facts in the form of triples, providing clearer insights into content factuality than traditional approaches. Through analyzing factual consistency across multiple LLM responses, we compute fine-grained hallucination scores without requiring external resources or training data. Our evaluation demonstrates that FactSelfCheck performs competitively with leading sentence-level sampling-based methods while providing more detailed and interpretable insights. Most notably, our fact-level approach significantly improves hallucination correction, achieving a $35.5\%$ increase in factual content compared to the baseline, while sentence-level SelfCheckGPT yields only a $10.6\%$ improvement. The granular nature of our detection enables more precise identification and correction of hallucinated content. Additionally, we contribute FavaMultiSamples, a novel dataset that addresses a gap in the field by providing the research community with a second dataset for evaluating sampling-based methods.
@@ 1 Introduction
Large Language Models (LLMs) have gained significant attention from academia and industry recently. However, a major limitation of LLMs is their tendency to generate hallucinated information (Farquhar et al., 2024; Huang et al., 2025), posing significant challenges for applications where factual correctness is crucial, such as healthcare (Sallam, 2023). Although numerous methods have been proposed to reduce hallucinations (Zhang et al., 2023), it is not possible to eliminate them, and LLMs will constantly hallucinate (Lee, 2023; Xu et al., 2024). Therefore, there remains a critical need for reliable hallucination detection in LLM responses, particularly for long-form text generation tasks where complexity and information density increase the risk of factual errors. Effective detection enables system interventions by either preventing the transmission of hallucinated content to users or facilitating its correction (Zhang et al., 2023).
Previous approaches to hallucination detection have primarily focused on classifying hallucinations at either the passage or sentence level (Huang et al., 2025). While valuable, these approaches are limited in their granularity and interpretability, as they do not provide detailed information about specific hallucinated facts or clear insights into what exactly is wrong. This limitation becomes particularly evident in long-form text generation. To address this limitation, we propose a novel method for hallucination detection that operates at the fact level, offering finer-grained and more interpretable analysis. In our approach, we define a fact as a triple consisting of a head, relation, and tail – a standard representation in knowledge graphs (e.g., (Robert Smithmember ofThe Cure)) (Hamilton et al., 2017). Our method provides more precise, actionable, and interpretable information by computing hallucination scores for individual facts than traditional passage-level or sentence-level classification approaches. This structured representation, through knowledge graphs, provides enables straightforward interpretation and verification (Pan et al., 2024).
@@
Our granular approach is motivated by two key observations. First, a single sentence can contain multiple facts, with the number of facts varying significantly across sentences, contexts, and domains. This variability makes it challenging to identify hallucinated aspects of generated output precisely when using sentence-level detection. Second, false information can be dispersed throughout a text, as a single fact may appear across multiple sentences. That can mislead the sentence-level detection, as the factuality of a sentence is dependent on the previous sentences (Zhang et al., 2024a). Fine-grained fact-level detection provides a more precise understanding of text factuality than a sentence analysis. It enables better assessment of content reliability and, as we show later, more effective factuality correction.
@@
We propose FactSelfCheck, a black-box method for fact-level hallucination detection, meaning it does not require access to the model’s internal parameters. This design choice makes our approach universally applicable across any LLM, including closed models, like GPT (OpenAI et al., 2024). Following a sampling-based detection paradigm, introduced by Manakul et al. (2023), our method utilizes multiple response generations and analyzes the factual consistency of extracted facts across these samples. This paradigm is based on the phenomenon that factual information remains largely consistent across different generations, while hallucinated content tends to vary or contradict itself between samples (Manakul et al., 2023; Wang et al., 2023). This way, we can effectively identify hallucinated facts without relying on external resources (zero-resource) or access to the model’s internal parameters (black-box). Our zero-resource, non-parametric approach requires no external knowledge bases or training data, making it broadly applicable across domains. The FactSelfCheck pipeline consists of three main steps: knowledge graph extraction, which extracts sets of facts from the initial response and samples; fact-level hallucination scoring; and calculating sentence-level scores by aggregating fact-level scores.
@@
We evaluated our method using the WikiBio GPT-3 Hallucination Dataset (Manakul et al., 2023) and the FavaMultiSamples dataset. WikiBio was the only existing dataset for evaluating sampling-based hallucination detection methods. To address this significant gap, we developed FavaMultiSamples, providing an additional evaluation benchmark. We performed both sentence-level and fact-level evaluation, with sentence-level scores aggregated from fact-level scores. Our approach achieves performance comparable to leading sampling-based methods at the sentence level while providing more detailed information about hallucinations. We also demonstrate effective fact-level hallucination detection and show that our fact-level approach significantly improves hallucination correction. Compared to a baseline, providing incorrect facts to the correction method leads to a $35.5\%$ increase in factual content, while passing incorrect sentences leads to an $10.6\%$ increase.
@@
    The novel zero-resource black-box sampling-based method for fact-level hallucination detection – FactSelfCheck, designed for long-form text generation. It enables fine-grained hallucination detection in LLM responses without requiring training data or external resources, as it is both non-parametric and zero-resource. We propose two effective approaches for measuring factual consistency across multiple samples: FactSelfCheck-KG using knowledge graph comparisons and FactSelfCheck-Text using direct text comparison.
@@ 2 Related work
Xu et al. (2024) have proven that hallucinations are inevitable in LLMs. As LLMs are powerful tools, many recent studies have been conducted regarding hallucination mitigation and detection (Zhang et al., 2023; Huang et al., 2025). The detection methods can be divided into two groups: white-box and black-box.
@@
Black-box approaches operate without access to the model’s internal states and aim to detect hallucinations based solely on the text generated by LLMs. Some of these methods use external resources to collect evidence (Min et al., 2023; Chern et al., 2023). Others leverage LLMs to detect hallucinations like CoVe (Dhuliawala et al., 2024), which utilizes the chain-of-thought paradigm for detection. Another category is sampling-based methods, such as SelfCheckGPT (Manakul et al., 2023), which evaluate factuality by generating multiple responses (stochastic samples) and assessing consistency between the original response and these samples. The paradigm of utilizing LLM to check its own responses was widely studied and adopted in many works (Kadavath et al., 2022; Lin et al., 2024; Ferraz et al., 2024; Zhang et al., 2024b; Miao et al., 2023). Many of these approaches employ a multi-step decomposition strategy to break down the complex task of hallucination detection into more manageable subtasks, a methodology we also adopt in our approach.
The most popular approach is to classify hallucinations at sentence-level or passage-level (Huang et al., 2025). Few methods have been specifically designed to detect hallucinations at the fact level, i.e. where facts are defined as triples. GraphEval (Sansford et al., 2024) generates a KG from LLM output and compares it with the context provided in the LLM input. FactAlign (Rashad et al., 2024) builds KGs from LLM output and source text, then compares them after performing entity alignment, a technique that pairs the same entity in different KGs. Knowledge-centric detection (Hu et al., 2024) similarly extracts knowledge triplets for fine-grained detection. All these methods require external context as reference, while our method is designed to work without any external knowledge sources by leveraging sampling-based consistency analysis. When evaluating on well-established sources like Wikipedia, methods with access to source materials may achieve superior performance (Manakul et al., 2023). However, our zero-resource approach offers broader applicability, as it can be applied to any task without requiring external knowledge sources.
Most similar to our approach is GCA (Fang et al., 2024), which constructs KGs from the response and samples and then compares them by aggregating multiple scores. However, GCA has significant methodological concerns – they tuned 6 hyperparameters directly on the evaluation set (the only available split in the WikiBio GPT-3 hallucination dataset (Manakul et al., 2023)), making it methodologically problematic. Due to these concerns, we cannot provide a fair quantitative comparison with GCA. In contrast, FactSelfCheck is truly zero-shot and parameter-free, requiring no parameter tuning. We achieved that by designing a constrained KG extraction that works consistently across multiple generation samples, rather than using freeform extraction like GCA.
@@ 3 Method
We propose FactSelfCheck, a black-box sampling-based method for fact-level hallucination detection, as illustrated in Figure 1. Our method is specifically designed for long-form text generation scenarios, where passages typically contain several sentences with complex factual information.
@@ 3.3 Knowledge Graph Extraction
We adopt an approach that decomposes the knowledge graph extraction task into simpler subtasks, similarly to Edge et al. (2024). This process involves three primary steps: extracting entities, identifying relations, and formulating facts, which are implemented as a sequence of LLM prompts.
@@ 3.4.1 FactSelfCheck-KG
where $\mathcal{H}_{\text{fact}}(f)$ is the hallucination score for fact $f$, and $\mathbb{I}\{f\in KG_{s_{n}}\}$ is an indicator function that equals $1$ if fact $f$ appears in $KG_{s_{n}}$ and 0 otherwise. The higher the $\mathcal{H}_{\text{fact}}$ value, the higher the plausibility of hallucination.
@@
[…]
@@ 3.5 Sentence-Level Hallucination Score
While detecting hallucinations at the fact level offers fine-grained insights, there are scenarios where sentence-level detection is necessary, such as for comparison with existing baselines. To achieve this, we aggregate fact-level scores to compute sentence-level scores $\mathcal{H}_{\text{sentence}}(u)$. This aggregation bridges the gap between atomic fact-level judgments and coarser sentence-level evaluations.
@@
where $u$ represents a single sentence, and $U$ is the set of sentences of the response $p$. The aggregation function Agg defines how the factuality of individual facts determines the factuality of the sentence. We employ two distinct aggregation strategies: meanmean and maxmax. The meanmean function computes the average hallucination score of all facts within a sentence, providing a smoothed measure of the overall factual density. This is useful for assessing general content quality where partial correctness matters. In contrast, the maxmax function identifies the most severe hallucination within the sentence, based on the intuition that even a single hallucinated fact is sufficient to identify the sentence as hallucinated.
@@ 4 Experimental Setup
In this section, we describe the experimental setup, including the used data, implementation details, and aspects we investigated. Additionally, we conducted a fact-level evaluation to provide a more comprehensive analysis. The detailed description of it is provided in Appendix C.2.
@@ 4.1 Evaluation Data
Since the method is designed for detecting hallucinations in long generated passages, finding appropriate datasets was challenging. We evaluated our method on two datasets. The first one is the WikiBio GPT-3 Hallucination Dataset (Manakul et al., 2023) 444huggingface.co/datasets/potsawee/wiki_bio_gpt3_hallucination (cc-by-sa-3.0 license) (later referred to as WikiBio). To the best of our knowledge, this was the only dataset specifically designed for evaluating sampling-based hallucination detection methods, representing a significant research gap. To address this, we developed FavaMultiSamples, a novel dataset specifically designed for evaluating methods that analyze multiple samples, providing researchers with an additional benchmark for robust evaluation. We built it upon the FAVA dataset (Mishra et al., 2024) and the detailed description of creation is provided in Appendix A. WikiBio covers generated biographical passages, while FavaMultiSamples includes diverse knowledge-intensive queries across various domains. Both datasets were annotated by humans. The dataset statistics are provided in Appendix B, and statistics related to KG extraction are detailed in Appendix C.4. Both datasets contain only test data, making it methodologically incorrect to tune parameters, including classification thresholds, on these datasets.
Sentence-level
While the datasets focus on sentence-level evaluation, our approach provides more fine-grained insights through fact-level analysis. To ensure a meaningful comparison with SelfCheckGPT, we evaluated these levels using aggregation approaches (see Section 3.5). For WikiBio, we followed the protocol established by Manakul et al. (2023), merging the labels major-inaccurate and minor-inaccurate into a single hallucination class.
Fact-level
To enable fact-level evaluation, we annotated facts from the extracted response knowledge graph ($KG_{p}$) using the LLM-as-judge approach (Zheng et al., 2023). For each fact, we provided the external biography from Wikipedia as a reference. The LLM-as-judge annotated whether each fact is supported by the reference. We utilized GPT-4o for this task, to ensure high quality of the annotations. As a result, we obtained $5488$ binary annotated facts.
The reliability of this approach is high, and commonly used in the literature (Thakur et al., 2025). The LLM-as-judge approach has been specifically tested for annotating hallucinations in LLM output given a knowledge source by Janiak et al. (2025) and has demonstrated high alignment with human annotations, outperforming other automated evaluation methods such as ROUGE (Lin, 2004).
@@ 4.2 Baseline Models
We compared our method against several key baselines. The RandomSentence baseline predicts a random class for each sentence with equal probability. The RandomFact baseline predicts random scores for each fact and aggregates them to obtain sentence-level scores. The probability-based baselines, proposed by Manakul et al. (2023), use -logpp and $\mathcal{H}$ to measure likelihood and entropy of each token respectively and aggregate the scores using Mean or Max functions to obtain sentence-level scores.
We include the two best-performing variants of SelfCheckGPT: Prompt and NLI. We implemented the Prompt variant using the same LLM employed in our method, namely Llama-3.1-70B-Instruct, whereas the original method utilized an unspecified release of GPT-3.5-turbo. For the NLI variant, we used the same model as was used in the original paper555huggingface.co/potsawee/deberta-v3-large-mnli (Apache license 2.0). Finally, the AttentionScore leverages attention maps (Sriramanan et al., 2024), which is, to the best of our knowledge, the only unsupervised internal state-based method. Its unsupervised nature was crucial since the available datasets do not provide training data. We adapted the AttentionScore for sentence-level detection by implementing two variants: (1) Absolute, which analyzes the complete attention map from the input start to the end of the given sentence, and (2) Relative, which analyzes on the attention map between the start and end of the given sentence.
@@
As discussed in Section 2, fact-level methods like GraphEval and FactAlign require external knowledge source as reference, while GCA tuned hyperparameters on test data. These design differences make quantitative comparison with our zero-resource approach methodologically inappropriate.
@@ 4.3 Implementation Details
We employed the Llama-3.1-70B-Instruct model (Grattafiori et al., 2024) as the LLM in all steps of our method and for baselines requiring access to LLM. We hosted it locally using vLLM (Kwon et al., 2023) on a server with 2 x Nvidia H100 94GB. For annotation of facts and hallucination correction, we utilized GPT-4o (OpenAI et al., 2024), as motivated in Sections 4.1 and 4.6. We set the LLM’s temperature to $0.0$ for all calls, except during hallucination correction (see Section 4.6), where we set it to $0.5$. The used prompts are available in the code repository. We implemented the methods and experiments using LangChain (Chase, 2022), Hugging Face Transformers (Wolf et al., 2020) and Hugging Face Datasets (Lhoest et al., 2021). All pipeline steps were defined using DVC (Kuprieiev et al., 2025) to facilitate reproducibility.
To determine whether a language model’s response was ’yes’ or ’no’ in Equation 9, we parsed the text into individual words and verified the presence of the words ’yes’ or ’no’. If either word was detected, the response was excluded from the averaging process in Equations 8 and 10. Moreover, our pipeline is vulnerable to not detecting facts in short, uninformative sentences; for these cases, we set the score to $0.5$. The analysis of the number of such sentences is provided in Appendix C.4.
4.4 Sentence-Level Detection